OpenStudy (anonymous):

Suppose Q and R are independent events. Find P(Q and R) if P(Q)=4/7 and P(R)=1/2

OpenStudy (anonymous):

hi Welcome to Openstudy! I believe \(P(Q\ and\ R)=P(Q) \times P(R)\) for independent events

OpenStudy (anonymous):

Apply this formula by substituting your given events: P(Q)=4/7 & P(R)=1/2